# Embedding in Recommender Systems: A Survey

### v2

@@ 6.5. Future Direction
Dynamic Codebook Quantization In online recommender systems, accommodating changing user preferences is crucial. Accurate quantization is essential for domain features that exert a significant impact on user preferences, like user interaction features. However, less impactful domain features like user location may need smaller codebooks. Exploring methods for capturing changes in user preferences and dynamically adjusting the codebook size for each domain feature accordingly is a promising research direction that warrants attention. Such research could considerably enhance the efficiency and effectiveness of online recommender systems.
7. Applications of Embedding in Deep Recommendation System
7.1. Embedding-based method for CTR tasks
With the advancement of various Embedding techniques, the capacity for expressive embedding has seen significant improvements. Embedding can now encode a wide range of features, rendering it a highly valuable technique. Moreover, mastering large-capacity embedding presents a formidable challenge. Consequently, it is advantageous to separate the training of embeddings from that of deep recommendation models. These attributes have made embedding pre-training an increasingly favored approach. Specifically, the typical procedure involves initially learning the dense vectors during the pre-training phase, followed by the incorporation of the well-trained embedding layer into the deep model for training the remaining parts of the model. Below are some practical examples of embedding pre-training in the context of industry applications, particularly in the domain of click-through rate (CTR) prediction.
Embedding pre-training with machine learning methods: The Factorisation-machine supported neural network (FNN) (Zhang et al., 2016b) represents a classic feed-forward neural network, with its foundational layer being an embedding layer initialized using parameters pre-trained on the factorization machine (FM) (Rendle, 2012b). Additionally, FNN has experimented with training the embedding layer through methods like Boltzmann machines(RBM) (Hinton, 2012) or denoising auto-encoders (DAE)  (Bengio et al., 2013). However, experiments in the realm of click-through prediction tasks demonstrate that the FNN model outperforms alternative techniques, including logistic regression, FM, RBM, or DAE-based embedding models. Facebook (He et al., 2014), on the other hand, employs a logistic regression model, with its input encoded via embeddings generated from a gradient-boosting decision tree, to accomplish ad click predictions. In this approach, the classification tree components function similarly to individual embedding layers that can be independently trained.
Embedding pre-training with deep neural network methods: Beyond initializing embeddings with machine learning models, the industry has achieved significant performance via end-to-end deep neural network (DNN) methods. For instance, the word2vec model (Mikolov et al., 2013b), introduced by Google utils a single-hidden layer neural network to pre-train the word embedding. Furthermore, Alibaba (Wang et al., 2018a) delves into graph embeddings through the Deepwalk (Perozzi et al., 2014) technique, facilitating large-scale e-commerce recommendations. Addressing the challenge of cold-start items lacking historical interaction data, Alibaba initializes cold-start item embeddings using side information from neighboring similar items with substantial user interaction. Beyond the two-tower architecture, attention mechanisms and transformer encoders present alternative architectures for embedding pre-training. For example, in Fig. 7, Alibaba (Zhou et al., 2018) combines the embedding layer and the well-designed MLP layer to make CTR prediction. In this method, DIN uses the attention mechanism to calculate the attention scores between the candidate ad and the goods the users bought.
Figure 7. The Architecture of DIN
While pre-training embeddings can enhance industrial deep learning models, it’s crucial to recognize that neural networks often require frequent or real-time training to adapt to the latest positive examples, especially in CTR prediction scenarios. This demand places a significant strain on computation resources and time. Therefore, it becomes essential to research and explore various training frequencies and strategies to strike a balance between training costs and model performance.
7.2. Embedding-based method for retrieval tasks
Modern recommender systems consist of three core components: the recall layer, ranking model, and re-ranking model. Due to computational limitations, real-world recommendation systems can not consider all items for ranking. Instead, the recall layer narrows down the selection to a smaller set of items, facilitating real-time recommendations. The re-ranking model, positioned closest to users, identifies relationships among ranked items from the ranking model and rearranges the candidate item sequence. Concurrently, the recall layer significantly impacts recommendation accuracy by reducing the candidate pool compared to the full set. However, if irrelevant items are included, it could degrade recommendation quality, making the recall layer a notable performance bottleneck in real-time systems. This section delves into embedding applications within the recall layer, encompassing single-tower and two-tower models.
Single-tower Model A prominent example of a single-tower recall model is the YouTube DNN video recall model (Covington et al., 2016). This model involves pre-learned embeddings for videos and search tokens. The recall process computes the average of watched video and search token embeddings, generating the user’s historical embedding. Furthermore, the user’s profile and contextual features combine with the historical embedding, forming input for an MLP. This input is condensed into a lower-dimensional space, yielding the final user embedding, known as the recall vector. During training, the softmax function gauges user-class probabilities and minimizes cross-entropy against the relevant label. To mitigate class imbalance and minimize the vast number of classes, negative sampling akin to Word2vec (Mikolov et al., 2013a) is employed. For recall, the nearest neighbor algorithm, e.g., Faiss (Johnson et al., 2019), retrieves $K$ video embeddings most akin to the user embedding.
Two-tower Model The single-tower model combines the user and item features and inputs them into an MLP to learn the user embedding for recall. However, this approach incurs a high computational cost for online computation. The two-tower model utilizes two separate neural networks to generate the user and item embeddings and then computes their similarity using the inner product. The two-tower model offers the advantage of being able to calculate the item embedding offline and only requires a calculation of the user embedding and similarity online. Additionally, the inner product operation in the two-tower model is more efficient compared to the softmax in the single-tower model, making it suitable for building recommendation systems with high real-time performance. The DSSM (Huang et al., 2013), introduced by Microsoft for computing semantic similarity, is one of the earliest two-tower models. It inputs a query and document and employs an MLP to learn the embeddings. The cosine similarity of the embeddings represents the semantic similarity between the query and the document, and the DSSM can be adapted for use as a recommendation model by replacing the query and document with user and item.
Both the DSSM and YouTube DNN Video Recall Models create negative samples through random sampling, but this approach is prone to generating false negative examples. This limitation could impact the performance of the recommendation system. The MOBIUS (Fan et al., 2019a), a two-tower model was proposed by Baidu to address the limitations of random sampling in creating negative samples. This model employs a teacher network to calculate the relevance score of the samples and selects those with low relevance scores as negative samples, thus providing a more effective strategy for mining difficult negative samples. The EBR model (Huang et al., 2020) proposed by Facebook uses the simplest two-tower structure, which uses the same hard negative sample mining strategy as MOBIUS. Further, EBR designs an embedding ensemble, which trains multiple models using different negative samples and integrates the output of the models using weighted or concatenated embedding. This design allows the model to effectively capture the diverse preferences and behaviors of users, leading to improved recommendation accuracy. EBR also finds that merging location features and social relationships into embedding has a significant impact on recommendation performance, which can also be used in subsequent studies.
Although the above-mentioned two-tower model has excellent processing speed, the intersection of user and item features in the two-tower model only occurs when the similarity between user embedding and item embedding is calculated, leading to a loss of information in the embeddings. As a result, this approach to decoupling the user and item in the two-tower model can greatly improve the speed of the recommendation system, but may also limit its accuracy. To perform feature intersection in the two-tower model, the DAT (Yu et al., 2021a) proposed by Meituan concatenates two augmented vectors onto the user and item feature vectors as the input to the two towers. The enhancement vectors are designed to capture the positive interactions of the other tower. Specifically, DAT calculates the MSE between the enhancement vector and the embedding of the other tower as the loss of feature interaction learning when the training sample is positive. When the training sample is negative, this part of the loss is ignored. To alleviate the sample category imbalance problem in industrial scenarios, DAT proposes category alignment loss to constrain the distance between the main category covariance matrix and other category covariance matrices.
Multi-interest Embedding Model Capturing multiple interests of users is one of the keys to recall, which prevents the recommendation system from being dominated by a single interest of users. The MIND model (Li et al., 2019), applied at Tmall, alleviates the limitation of recall ability for niche content in the recommender system by incorporating a multi-interest extraction layer to capture users’ potential multiple interests in the user-item click sequence. It designs a capsule network-based multi-interest extraction layer in the recall layer, which decomposes the embedding interactions of sequential items into multiple embeddings. The SDM (Lv et al., 2019) applied to Taobao sequence recall utilizes a self-attention mechanism to capture multiple interests of users within a session, and a gated neural network to fuse the long-term and short-term interests of users. Specifically, SDM defines a session based on the time window of the interaction (e.g., an interaction belongs to the same session if the interaction difference is below a certain threshold, with a maximum session length limit). The most recent session is considered a short-term behavior, while earlier sessions are considered long-term behaviors. Since the shorter sequence of items visited by short-term behavior allows for designing a more complex model, SDM uses LSTM and self-attention mechanism to learn its embedding, while long-term behavior applies attention mechanism and simple fully connected layer to learn the embedding, and finally, the recall vector is obtained by fusing long-term and short-term interests through the gated neural net. The Airbnb listing recommendation system (Grbovic and Cheng, 2018) generates embeddings of users and listings for recall or ranking based on user click sessions and booking sessions. It utilizes a skip-gram model to encode the sequence of clicked listings to capture users’ short-term interests and make real-time, personalized recommendations within a session. Meanwhile, Airbnb wants to capture users’ long-term interests from their historical booking session sequences, but due to the sparse booking behavior of users, the recommendation system faces a serious cold-start problem, which makes it difficult to learn meaningful embeddings. To alleviate this problem, Airbnb proposes an embedding learning method based on similar users and similar listings. Specifically, Airbnb clusters user and item separately and treats user type as user id, generates a booking session consisting of listing type, and learns the embedding of user type and item type, this idea is similar to group recommendation.
Approximate Near Neighbor Embedding Search The above-mentioned model aims to train the user embeddings for recall, and after getting the user embeddings use an approximate near neighbor embedding search algorithm to recall the item ID. Most of the existing work uses Faiss (Johnson et al., 2019), Facebook’s open-source similarity search library that can compute efficiently on GPU, to complete the recall task. Milvus (Wang et al., 2021d) integrates the functionality of several vector similarity search libraries (e.g. Faiss (Johnson et al., 2019), Microsoft SPTAG (Chen et al., 2018)) to develop a high-performance vector similarity search system for distributed systems and dynamic data.
7.3. Empowering Recommendation Systems with Enhanced Embeddings Using Large Language Models
Recently, there has been a notable surge of interest in the incorporation of Large Language Models (LLMs) into embedding-based recommendation systems (Wu et al., 2023). For the enhancement of embeddings in recommendation systems with LLMs, two prevalent strategies emerge:
@@ 7.3. Empowering Recommendation Systems with Enhanced Embeddings Using Large Language Models
Direct Inference of Embeddings Utilizing LLMs. In this strategy, using LLMs, we can extract embedding vectors from textual descriptions of users, items, and user-item interactions. This is particularly effective in scenarios with limited data for user-item interactions. For example, ChatRec (Gao et al., 2023) generates embeddings through prompting a LLM with prompts generated by a multi-input prompt constructor module. The module forms a natural language paragraph that captures user intent and item details by considering various inputs, such as user-item history interactions, user profiles, specific user queries, and dialog history. GeneRec (Wang et al., 2023) uses user profiles, historical feedback, and collected Web data (e.g., common facts and knowledge from Wikipedia) to create personalized embeddings through LLMs. Specifically, it can generate personalized embeddings by analyzing user instructions (e.g., fusing images, audio, and natural languages) and feedback (e.g., clicks), resulting in tailored content like landscape micro-videos in the chosen style.
Fine-tuning Embeddings with LLMs. LLMs can leverage user-item interactions through self-supervised fine-tuning, enhancing the quality of embeddings and subsequently boosting the overall performance of the system. This approach treats user or item IDs as distinct untrained tokens and optimizes them with a self-supervised fine-tuning process, such as generative retrieval tasks (Wu et al., 2022). For example, in the case of PEPLER (Li et al., 2023b), it successfully embedded IDs into a semantic space for better embedding explanation ability. PEPLER utilizes a two-stage approach for fine-tuning embedding vectors and model parameters. In the first stage, the task of generating recommendation explanations relies on a sequential fine-tuning procedure for next-token prediction, using a negative log-likelihood (NNL) loss. Initially, the ID embedding vectors are trained with the LLM frozen, followed by a fine-tuning procedure that updates both the LLM and the vector parameters. In the second stage, two tasks including next-token prediction by minimizing the NNL loss and rating prediction minimizing the mean square error, are both employed to improve the generated explanation along with recommendation performance. PPR (Wu et al., 2022) incorporates contrastive loss as auxiliary losses by employing two unique data augmentation techniques: prompt-level augmentation involves randomly masking elements in user embedding vectors, and behavior-level augmentation that masks some items in the user-item interaction sequence. This allows for the learning of personalized embedding vectors that harness the information from user profiles.
@@
LLMs can play an important role in recommendation systems by extracting features and enhancing embeddings through the incorporation of the rich semantic information embedded in the pretrained and fine-tuned LLMs. LLM-enhanced embeddings have demonstrated their potential in enhancing performance in applications like video recommendation (Liu et al., 2023a) and recommendation explanation generation (Zhou and Joachims, 2023); however, the challenge of effectively integrating LLMs with recommendation systems remains an open problem (Wu et al., 2023). For example, LLMs demand substantial computational resources and time, which may be unfeasible in resource-constrained and real-time recommendation scenarios. Additionally, LLMs have exhibited biases (Liu et al., 2023b), like gender bias (Thakur, 2023), which can potentially lead to untrustworthy recommendations.
Although there have been notable improvements in the integration of LLMs with recommendation systems, research on embedding techniques is still in its nascent phase. As far as we know, there have been no comprehensive attempts to summarize and investigate the role of embedding approaches in recommendation systems that depend on large language models. This subsection, therefore, presents a preliminary groundwork. We anticipate that there will be a significant focus on exploring embeddings in the context of LLM-based recommendation systems.

### v3

@@ 7.4. Surveys and Future Directions
Future Direction 2: Dynamic Codebook Quantization. In online recommender systems, accommodating changing user preferences is crucial. Accurate quantization is essential for domain features that exert a significant impact on user preferences, like user interaction features. However, less impactful domain features like user location may need smaller codebooks. Exploring methods for capturing changes in user preferences and dynamically adjusting the codebook size for each domain feature accordingly is a promising research direction that warrants attention. Such in-depth research could considerably enhance the operational efficiency and overall effectiveness of online recommender systems.
8. LLM-Driven Embedding Enhancement
Recent years have witnessed remarkable advancements in the field of large models, particularly Large Language Models (LLMs), which have demonstrated transformative potential across various domains. These models, with their strong language understanding and reasoning capabilities, are now increasingly being leveraged within recommender systems to overcome the limitations of traditional embedding techniques and further elevate the quality of recommendations. This chapter introduces the paradigm of ”Large Model Enhanced Modeling” in recommender systems, where the power of these models is harnessed to refine and augment the embedding layer, leading to significant improvements in capturing user preferences and item characteristics. This section will explore three key areas where large models are making a substantial impact on embedding techniques: (1) LLM embedding as the direct semantic supplement, and (2) LLM embedding as guidance. By delving into these aspects, this section aims to provide a comprehensive understanding of how large models complement and extend the capabilities of existing centralized embedding methods discussed in this survey, including matrix factorization, sequential modeling, and graph embeddings, as well as techniques addressing scalability like AutoML, hashing, and quantization. Rather than replacing these established methods, large models offer a powerful means of enhancing their effectiveness. An overview of these enhancement strategies is provided in Table 8.
Table 8. LLM-Driven Embedding Enhancement
LLM-Driven Embedding Enhancement LLM Embedding as Semantic Supplement (§8.1) Direct Inference: ChatRec (Gao et al., 2023), GeneRec (Wang et al., 2023a), TedRec (Xu et al., 2024), LRD (Yang et al., 2024b), SemSR (Narwariya et al., 2025)
With Trainable Modules: LLM-ESR (Liu et al., 2024b), AlphaRec (Sheng et al., 2024), Laser (Lin et al., 2025), SRA-CL (Cui et al., 2025), SAID (Hu et al., 2024)
Fine-tuning LLM/Embeddings: PEPLER (Li et al., 2023c), PPR (Wu et al., 2022), LLM-CF (Sun et al., 2024), TPAD (Su et al., 2025), LLMEmb (Liu et al., 2025b), SeRALM (Ren et al., 2024a), LLM2Rec (He et al., 2025)
LLM Embedding as Guidance (§8.2) Self-Guidance: LLM4SBR (Qiao et al., 2024b), DaRec (Yang et al., 2025)
Cross-Embedding Guidance: Jia et al. (Jia et al., 2025), PAD (Wang et al., 2025a), LLM4CDSR (Liu et al., 2025d), EIMF (Qiao et al., 2024a), LGMcRec (Meng et al., 2025)
8.1. LLM Embedding as Semantic Supplement
Natural languages contain various information about users and items, but it is hard to effectively add textual features into recommender systems. Recently, there has been a notable surge of interest in the incorporation of Large Language Model embeddings into recommendation systems to supplement semantic information (Wu et al., 2023; Liu et al., 2025c; Xiao et al., 2025). In the recommendation task, given the interaction sequence of a given user $u$ can be denoted as $\mathcal{\boldsymbol{S}}^{(u)}=[v_{1}^{(u)},v_{2}^{(u)},\dots,v_{|\mathcal{S}^{(u)}|}^{(u)}]$, the aim is to predict the next item the users would like to select, which can be formulated as:
(25) $\arg\max_{v_{j}\in\mathcal{V}}P(v_{N+1}=v_{j}|\mathcal{\boldsymbol{S}}),$
where $v_{j}\in\mathcal{V}$ is the selected item in the item set, $N$ is the number of items that the user interacts. In the LLM-driven embedding enhancement strategies, the embedding procedure of the item sequences is necessary, which can be written as $\text{LLM}_{emb}(\cdot)$. This function usually denotes that the last hidden state of the LLM is regarded as the embedding of a given item sequence. However, the item ids are labels that LLMs are hard to understand, so a prompt $T_{i}=[I,A_{1},A_{2},\dots,A_{K}]$ should also be constructed for the $i$-th item, where $I$ is the instruction for LLM generation, and $A_{k}(k\in[1,K])$ is the attribute of the item. Then the LLM-driven embedding $\boldsymbol{e}_{i}^{\text{LLM}}$ can be generated:
(26) $\boldsymbol{e}_{i}^{\text{LLM}}=\text{LLM}(T_{i}),\ \ i\in[1,N].$
As shown in Figure 8, these LLM embeddings can be utilized for direct inference, and they can be fine-tuned for capturing more fine-grained semantic relationships. In this subsection, we will discuss the two categories in detail.
Figure 8. Categories of LLM Embeddings as Semantic Supplement. (a) Generate embeddings using frozen LLMs and convey them to recommendation models after dimensional reduction. (b) Employ a trainable adapter to relieve the information loss of LLM embeddings. (c) Fine-tune LLMs or the embeddings to enhance the semantic information for recommendation tasks.
@@ 8.1. LLM Embedding as Semantic Supplement
Direct Inference of Embeddings Utilizing LLMs. In this strategy, using LLMs, we can extract embedding vectors from textual descriptions of users, items, and user-item interactions. This is particularly effective in scenarios with limited data for user-item interactions. For example, ChatRec  (Gao et al., 2023) generates embeddings through prompting a LLM with prompts generated by a multi-input prompt constructor module. The module forms a natural language paragraph that captures user intent and item details by considering various inputs, such as user-item history interactions, user profiles, specific user queries, and dialog history. GeneRec  (Wang et al., 2023a) uses user profiles, historical feedback, and collected Web data to create personalized embeddings through LLMs. Specifically, it can generate personalized embeddings by analyzing user instructions and feedback, resulting in tailored content like landscape micro-videos in the chosen style. TedRec  (Xu et al., 2024) and LRD  (Yang et al., 2024b) utilize frozen language models to directly generate representations for embeddings of semantic information and IDs. SemSR  (Narwariya et al., 2025) generates the semantics embedding from the LLMs and integrates it into a trainable end-to-end Semantic sequential recommender.
However, directly integrating such embeddings might lead to information loss  (Liu et al., 2025c). Therefore, recently, more research has focused on enhancing the embeddings from LLMs by adding trainable modules to tailor the embeddings to LLMs more effectively. One representative work is LLM-ESR  (Liu et al., 2024b), where a trainable adapter module is designed for effective original semantics retention in both the trainable collaborative embedding layer and trainable cross attention. In AlphaRec  (Sheng et al., 2024), an MLP serves as a bridge between the language representations and ID representations. SRA-CL  (Cui et al., 2025) utilizes the sequence encoder for user embeddings from LLMs to enhance the information representation capability of contrastive samples. Laser  (Lin et al., 2025) employs an adapter to enhance features by leveraging the user and item embeddings derived from the LLM. In addition, trainable encoder modules can also be applied to generating global interaction embeddings and learning user preferences in the cross-domain recommendation  (Liu et al., 2025d). SAID  (Hu et al., 2024) semantically aligned the item embeddings for sequential recommenders using a projector to derive an embedding vector, which is utilized to preserve fine-grained semantic information. Uniquely, the projector is aligned within the LLM for better embedding generation, followed by the training together with the recommenders.
Fine-tuning LLM-Driven Embeddings. Embeddings can be enhanced not only by adding a trainable module to learn more effective representations, but also by fine-tuning embedding vectors directly or by fine-tuning LLM itself to adapt the embedding generation, thereby enhancing the quality of embeddings and boosting the overall performance of the system. For example, in the case of PEPLER (Li et al., 2023c), it successfully embedded IDs into a semantic space for better embedding explanation ability. PEPLER utilizes a two-stage approach for fine-tuning embedding vectors and model parameters. PPR (Wu et al., 2022) learns personalized embedding vectors by employing prompt-level augmentation, which involves randomly masking elements in user embedding vectors, and behavior-level augmentation that masks some items in the user-item interaction sequence. LLM-CF  (Sun et al., 2024) designs a data mixture method to distill the world knowledge of LLMs into recommenders, also enhancing the representation ability of item embeddings.
Since the LLMs are the key source of embeddings’ semantic information, fine-tuning LLMs can also be effective in conveying more knowledge to the embeddings. For example, TPAD  (Su et al., 2025) fine-tunes the multimodal LLMs to generate better knowledge item embeddings, thus aligning both transitional patterns and knowledge patterns in one unified LLM more effectively. LLMEmb  (Liu et al., 2025b) employs a supervised contrastive fine-tuning approach to train the LLM together with the adapter module to learn effective information from LLM embeddings. Laser  (Zhang et al., 2024a) uses a bi-tuning method by inserting trainable virtual prefix and suffix tokens into the input sequence. These semantic tokens enhance the LLMs’ capability of generating high-quality embeddings. SeRALM  (Ren et al., 2024a) applies an alignment training to refine LLMs’ embedding generation according to the feedback from the sequential recommenders, achieving a cycling optimization paradigm. Moreover, the fine-tuning of embeddings can be combined with fine-tuning of LLMs in one framework to collaboratively learn better embedding utilization. For instance, LLM2Rec  (He et al., 2025) employs an innovative two-stage training framework that uses a trainable LLM to generate embedding and inherently enhances both semantic information and collaborative filtering signals rather than merely using the LLM as a feature extractor or training guide.
8.2. LLM Emebdding as Guidance
Unlike approaches that directly utilize LLM embeddings as input features, this enhancing strategy treats embeddings not as primary representations, but rather as high-level semantic guidance to steer the training process or facilitate parameter synthesis  (Liu et al., 2025c). By decoupling representation learning from downstream model adaptation, these methods aim to transfer and guide the learning process, as well as enhance the embedding’s expressive capability in recommendation models. The guiding mechanism of the LLM embeddings can be categorized into Self-Guidance and Cross-Embedding Guidance, which will be introduced in detail in this subsection.
Self-Guidance. LLM embeddings can be used as a guide to identify representative items of users. LLM4SBR  (Qiao et al., 2024b) and LLM-ESR  (Liu et al., 2024b) enhance the performance of recommendation models by utilizing LLM embeddings to identify user preferences and similar user behaviors, thereby refining the modeling of diverse behavior patterns and guiding the learning process of embedding utilization in recommendation models. DaRec  (Yang et al., 2025) proposes a plug-and-play disentanglement framework that separates LLM and collaborative embeddings into shared and task-specific components, then aligns the shared representations. LLMEmb  (Liu et al., 2025b) employs a supervised contrastive fine-tuning approach to facilitate self-guided training for the effective utilization of LLM embeddings, thereby alleviating the long-tail problems in recommendation.
Cross-Embedding Guidance. Basic embeddings from users and items, and the LLM embeddings of users and items, share potential semantic relationships. In addition, they contribute different semantic information for recommenders. Therefore, the LLM embeddings can also serve as guidance for learning basic embeddings. Jia et al.  (Jia et al., 2025) demonstrate a typical cross-embedding guidance paradigm, where the LLM embedding of users and items can be applied to refine the corresponding basic embeddings. PAD  (Wang et al., 2025a) constructs LLM embeddings of item titles and aligns them with the initial item ID embeddings. LLM4CDSR  (Liu et al., 2025d) employs the global interaction embeddings from LLMs and hierarchical LLMs profile embeddings to guide the learning of local user embeddings and item embeddings. EIMF  (Qiao et al., 2024a) constructs the embeddings of clustered user behaviors, which are also used to guide the learning of basic behavior embeddings. LGMcRec  (Meng et al., 2025) uses the LLM to generate user and item profile embeddings, which are applied to enhance the embeddings of user and item interaction embeddings learned from graph neural networks.
8.3. Surveys and Future Directions
@@
Survey: LLMs can play an important role in recommendation systems by extracting features and deriving high-quality embeddings through the incorporation of the rich semantic information embedded in the pretrained or fine-tuned LLMs. These embeddings can be enhanced through the specially designed and trained adapter modules or fine-tuned methods on the embeddings. LLM-enhanced embeddings have demonstrated their potential in enhancing performance in applications like video recommendation (Liu et al., 2023a) and recommendation explanation generation (Zhou and Joachims, 2023).
Future Direction 1: Computational Efficiency and Deployment. While LLMs show great promise for enhancing recommendation systems through embedding enhancement, the computational demands pose significant implementation barriers (Wu et al., 2023). These models require substantial computational resources for inference and fine-tuning, making them challenging to deploy in real-world recommendation scenarios where millisecond-level response times are often required (Zhang et al., 2025, 2024b). This challenge necessitates research into efficient embedding caching strategies, model compression techniques, and lightweight adapter architectures specifically optimized for recommendation contexts. Integration challenges also present significant hurdles in determining optimal strategies for combining LLM-derived semantic knowledge with traditional collaborative filtering signals, maintaining high recommendation quality while reducing computational overhead, and striking the right balance between personalization and generalization when utilizing LLM embeddings (Wu et al., 2023). Scalability remains a critical concern, particularly in handling large-scale user-item matrices and efficiently updating LLM-enhanced embeddings as new items and users are added to the system (Zhang et al., 2025).
Future Direction 2: Fairness and Privacy Enhancement. Another crucial concern relates to bias and fairness in LLM-enhanced recommendations. LLMs can inherit and amplify societal biases present in their training data, including gender, racial, and other demographic biases, which may lead to unfair or discriminatory recommendations (Liu et al., 2023b; Thakur, 2023). This calls for dedicated research into methods for detecting and mitigating bias in LLM-enhanced embeddings, developing fairness-aware fine-tuning approaches, and establishing transparent evaluation metrics for assessing recommendation fairness. Privacy considerations in processing user data through LLMs add another layer of complexity to these integration challenges (Liu et al., 2023b). Developing effective techniques for addressing cold-start problems while leveraging LLM knowledge is also essential. These challenges collectively present important research opportunities for making LLM-enhanced recommendation systems more practical and trustworthy for real-world applications.